OpenStudy (okdutchman7):

Please help. WILL MEDAL! Inverse functions @Hero @sleepyjess @jim_thompson5910 @robtobey @Erak

OpenStudy (okdutchman7):

Determine if \[f(x) = \frac{ 1 }{ 1-x }\] and \[g(x) = \frac{ x-1 }{ x }\] are inverse functions by computing their compositions.

OpenStudy (okdutchman7):

\[\frac{ 1 }{ 1-\frac{ (x-1) }{ x } }\]

OpenStudy (okdutchman7):

\[\frac{ 1 }{ \frac{ x-x+1 }{ x } }\]

OpenStudy (okdutchman7):

\[\frac{ x }{ 1 }\]

OpenStudy (okdutchman7):

x

jimthompson5910 (jim_thompson5910):

you've shown that f(g(x)) = x for all x in the domain

jimthompson5910 (jim_thompson5910):

now you need to show if g(f(x)) = x is true

OpenStudy (okdutchman7):

The answer on my answer sheet says "yes" but I don't know why something is yes or no.

OpenStudy (okdutchman7):

oh ok.

jimthompson5910 (jim_thompson5910):

The original question is "are f and g inverses of each other" the answer is "Yes" I'm assuming your teacher wants you to show why "yes" is the correct answer

OpenStudy (okdutchman7):

\[\frac{ \frac{ 1 }{ 1-x }-1 }{ \frac{ 1 }{ 1-x } }\]

OpenStudy (okdutchman7):

I don't know what would make the answer yes, though.

jimthompson5910 (jim_thompson5910):

simplify that last expression you wrote

OpenStudy (okdutchman7):

\[\frac{ \frac{ 1-(1-x) }{ 1-x } }{ \frac{ 1 }{ 1-x } }\]

OpenStudy (okdutchman7):

\[\frac{ x(1-x) }{ 1-x }\]

OpenStudy (okdutchman7):

x

OpenStudy (okdutchman7):

So I got x both ways. Is that significant?

jimthompson5910 (jim_thompson5910):

so because f(g(x)) = x g(f(x)) = x this means that f and g are inverses of each other

OpenStudy (okdutchman7):

So if I don't get x than they are not inverse?

jimthompson5910 (jim_thompson5910):

if you got f(g(x)) to be equal to something other than x, then f and g wouldn't be inverses of each other. or if you got g(f(x)) to be equal to something other than x, then f and g wouldn't be inverses of each other.

OpenStudy (okdutchman7):

So if I don't get x than they are not inverse?

jimthompson5910 (jim_thompson5910):

if f and g are inverses of each other, then f(g(x)) = x and g(f(x)) = x for all defined x values in the domain

jimthompson5910 (jim_thompson5910):

`So if I don't get x than they are not inverse?` yes

OpenStudy (okdutchman7):

Thank You!

jimthompson5910 (jim_thompson5910):

no problem